mike wants to make meatloaf. his recipe uses a total of 5 pounds of meat. if he uses a 3 to 1 ratio of beef to pork how much pork will he use?
Answer by checkley77(12844)

3+1=4 total shares.
5/4=1.25 value of each share.
3*1.25=3.75 for the amount of beef
1*1.25=1.25 pounds of pork
proof:
3/1=3.75/1.25
3*1.25=1*3.75
3.75=3.75